# Gamma/hadron discrimination through the analysis of the shower footprint at low energies

### v1

@@ IV.3 Dependence on the muon content of the shower
Figure 7: 2D histogram showing the method’s probability as a function of the number of muons in stations for proton showers with $E_{r}\sim 500\,\rm GeV$. Mean values are shown with error bars representing the standard deviation for both proton (red) and gamma (blue) showers in this reconstructed energy bin.
@@ V Conclusions
In this study, the gamma/hadron separation capability of a footprint-based method across various conditions, including the presence of atmospheric muons, low-energy proton shower noise, and differing shower core positions was evaluated.
The method demonstrates strong resilience against noise from low-energy proton showers and atmospheric muons, with negligible impact on the separation efficiency across the energy bins tested. The method consistently outperformed the HAWC and LHAASO experiments across all tested conditions.
It was observed that the performance shows minimal dependence on the position of the shower core, with only a slight degradation noted between events closer to and farther from the center of the array at the lowest reconstructed energy bin, where shower footprints are smaller. This suggests that the classification performance remains robust across different core positions.
Finally, although the presence of muons enhances the performance, it is still effective even in their absence. This demonstrates that the model is not solely dependent on muons but also leverages other footprint features for gamma/hadron separation.
Overall, the ViT-based approach shows promise in achieving efficient gamma/hadron separation under realistic conditions, making it a viable candidate for use in current and future gamma-ray observatories.
@@ Acknowledgments
We would like to express our gratitude to Antonio Bueno for his valuable discussions during the early stages of this work. This work has been financed by national funds through FCT - Fundação para a Ciência e a Tecnologia, I.P., under project PTDC/FIS-PAR/4320. B.S.G. (LIP/IST) is grateful for the financial support from the FCT PhD grant PRT/BD/151553/2021 under the IDPASC program (https://doi.org/10.54499/PRT/BD/151553/2021). A.G. is grateful for the financial support from the Spanish Ministry of Economy and Competitiveness (MINECO) under the project PID2019-104676GB-C32.

### v2

@@ IV.6 Dependence on the muon content of the shower
Figure 11: 2D histogram showing the method’s probability as a function of the number of muons in stations for proton showers with $E_{r}\sim 500\,\rm GeV$. It is important to note that the color bar considers only proton-induced events. Mean values are shown with error bars representing the standard deviation for both proton (red) and gamma (blue) showers in this reconstructed energy bin.
@@ V Conclusions
In this study, the gamma/hadron separation capability of a footprint-based method was evaluated under various realistic conditions. Factors such as the presence of atmospheric muons, low-energy proton shower noise, variations in shower core locations, zenith angles, and array fill factors were assessed.
It was proved that, although the presence of muons enhances the performance, the method is still effective even in their absence. This demonstrates that the model is not solely dependent on muons but also leverages other footprint features for gamma/hadron separation, which is crucial at the sub-TeV energy range where muons are scarce or even absent.
The method demonstrated strong resilience to noise from low-energy proton showers and atmospheric muons, with negligible impact on the separation efficiency across the energy bins tested. Under identical conditions, this approach suggests improved performance and noise resilience compared to other state-of-the-art gamma/hadron discriminators such as the ones designed at the HAWC gamma-ray observatory.
For inclined showers with zenith angles of $\sim 30^{\circ}$, the method proved to be effective with a similar classification performance observed for nearly vertical events. This highlights the robustness of the proposed approach, confirming its applicability to both near-vertical and inclined events.
With respect to dependence on the location of the shower core, the efficiency in discrimination for on-array events is stable, with minor variations in the average output probabilities at different core distances. However, the classification becomes more challenging for off-array events, with higher fluctuations for gamma-ray events. Nervertheless, the results suggest that the proposed method remains effective for core distances up to $180\,\rm m$, surpassing the array radius of $160\,\rm m$. This proves its robustness across various core positions, both within and near the boundaries of the array.
The investigation into different array fill factors revealed that the gamma/hadron separation efficiency is not significantly affected by the density of the detector array. Lower fill factors result in a slight reduction in classification efficiency, primarily due to the loss of granularity in the shower footprint. However, the proposed technique maintains strong separation capabilities even for a fill factor as low as $40\%$, indicating its adaptability to different detector configurations.
Overall, this footprint-based method showed exceptional background rejection across all tested scenarios, with a performance similar to that of state-of-the-art techniques in current gamma ray experiments. These results demonstrate that there is enough information in the shower footprint to perform an outstanding gamma/hadron separation, making it a viable candidate for use in current and future gamma-ray observatories, improving their sensitivities at sub-TeV energies.
@@ Acknowledgments
We would like to express our gratitude to Antonio Bueno for his valuable discussions during the early stages of this work. This work has been financed by national funds through FCT - Fundação para a Ciência e a Tecnologia, I.P., under project PTDC/FIS-PAR/4320. B.S.G. (LIP/IST) is grateful for the financial support from the FCT PhD grant PRT/BD/151553/2021 under the IDPASC program (https://doi.org/10.54499/PRT/BD/151553/2021). A.G. is grateful for the financial support from the Spanish Ministry of Economy and Competitiveness (MINECO) under the projects PID2019-104676GB-C32 and PID2023-147949NB-C53.
Appendix
V.1 Optimal gamma selection efficiency
Figure 12 presents the background rejection factor for a gamma efficiency selected to achieve the optimal Q-factor, as defined in Equation 1.
$Q=\frac{\varepsilon_{\gamma}}{\sqrt{\varepsilon_{p}}}$ (1)
Figure 12: The gamma (top) and proton (bottom) selection efficiency at different energies and noise sources, with the gamma selection efficiency set to achieve the optimal Q-factor.
V.2 Implementation of other gamma/hadron separators
Figure 13: The gamma (top) and proton (bottom) selection efficiency achieved by the HAWC gamma/hadron separators evaluated for showers without noise compared with HAWC at a lower altitude (represented by the gray dotted line) [58].
The principal HAWC gamma/hadron discriminators were adapted and incorporated into our simulation framework. Although HAWC typically calculates these variables using the charges at the PMT level, in this scenario, given the reduced size of the WCD, enhanced results were obtained by using the total charge from the three PMTs in the Mercedes WCD. The MC core was used for the calculation of these discrimanators.
The first variable introduced was LIC, which is the logarithmic transformation of the inverse of the compactness parameter, an empirical parameter originally developed by the Milagro Collaboration [59, 60]. These variables were computed as follows:
[…]
where the CxPE40 represents the charge detected by the WCD located more than $40\,\rm m$ from the shower core, which exhibits the highest effective charge, and N is the total number of active stations. In stations with muons, it is anticipated that larger signals will be observed, consequently high LIC values are expected for showers initiated by hadrons.
Other discriminators can be obtained using the expected LDF of gamma rays, such as the PINCness and the LDF $\chi^{2}$. The PINCness is a $\chi^{2}$-like measure of the smoothness of the charge footprint of the showers [59]. To achieve this, as detailed in Equation (3), it calculates the signal discrepancy between each WCD, $q_{i}$, and the average signal at stations equidistant from the shower core, $\left\langle\log_{10}\left(q_{i}\right)\right\rangle$, organising the stations into concentric rings with a width of $10\,\rm m$. The discrepancy is normalised using the expected charge error $\sigma$, which was calculated for each energy bin using the stations within the gamma-ray events of the training datasets. Subsequently, a $\chi^{2}$-like metric is obtained by dividing the sum by the degrees of freedom, which equates to the count of stations, denoted as N.
[…]
Alternatively, the LDF $\chi^{2}$ employs a modified Nishimura-Kamata-Greisen (NKG) function [58], as described in Equation (4), to estimate the logarithm of the signal of a station, $\log_{10}\left(q_{\text{LDF }}\right)$, at a certain distance from the shower core, $r_{\rm core}$. Following this, the LDF $\chi^{2}$ is computed according to Equation (5) utilizing HAWC’s Moliere radius: $124.21\,\rm m$.
$\log_{10}(\text{NKG})$ $=\log_{10}A+s\left(\log_{10}\left(\frac{r_{\rm core}}{124.21}\right)\right.$ (4)
$\left.+\log_{10}\left(1+\frac{r_{\rm core}}{124.21}\right)\right)$
$-3\log_{10}\left(\frac{r_{\rm core}}{124.21}\right)$
$-4.5\log_{10}\left(1+\frac{r_{\rm core}}{124.21}\right)$
$\chi^{2}=\frac{1}{N}\sum_{i=0}^{N}\frac{\left(\log_{10}\left(q_{i}\right)-\log_{10}\left(q_{\text{LDF }}\right)\right)^{2}}{\sigma^{2}}$ (5)